# Elements of World Knowledge (EWoK): A cognition-inspired framework for evaluating basic world knowledge in language models

### v1

@@ Acknowledgements
We thank everyone who contributed ideas on dataset design, especially Hayley Ross and Yuhan Zhang. We thank members of LINGO Lab at MIT for providing feedback on earlier versions of the work. This work was supported by the Language Mission of the MIT Quest for Intelligence.
@@ Appendix A Methods
A.1 Human data collection
@@
To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item. To control for quality of data, we excluded 59 participants whose inter-subject correlation (compared with average ratings on items from other participant who rated the same items as this subject) was below . Each EWoK-core-1.0 generated item (corresponding to 5 variable-populated versions) was split into four subparts: and presented in a Likert setting using the same prompt as that used for LLMs (Appendix A.4; adapted to presentation in a web browser with a free-form text-box for human input). Items were presented so that each participant only saw one of 5 variable-populated variants and one of the four possible sub-items . Therefore, participants provided sensibility judgments independently of any other subparts of the item, closely matching the conditions for LLM evaluation. Each participant saw a list of 57 items on average, depending on the domain (188 for agent properties and 27 for material properties).
A.2 Evaluated models
The full set of evaluated models are as follows: gpt2-xl (Radford et al., 2019), phi-1 (Gunasekar et al., 2023), phi-1.5, phi-2 (Li et al., 2023), gemma-2b, gemma-1.1-2b-it, gemma-7b, gemma-1.1-7b-it (Gemma et al., 2024), mpt-7b, mpt-7b-chat, mpt-30b, mpt-30b-chat (MosaicML, 2023), falcon-7b, falcon-7b-instruct, falcon-40b, falcon-40b-instruct (Almazrouei et al., 2023), mistral-7b-v0.1 (Jiang et al., 2023), mixtral-8x7b-v0.1 (Jiang et al., 2024), Meta-Llama-3-8B, and Meta-Llama-3-70B (AI@Meta, 2024). All models were accessed via HuggingFace transformers (Wolf et al., 2019), and all experiments were run on a x GB GPU cluster.
A.3 Text completion experiments
@@ A.4 Prompts
A.5 Mixed effects modeling

### v2

@@ Acknowledgements
We thank our action editor Emiel Krahmer and the anonymous reviewers for providing valuable feedback that helped improve this work. We thank everyone who contributed ideas on dataset design, especially Hayley Ross and Yuhan Zhang, members of LINGO Lab at MIT for providing feedback on earlier versions of the work, and Yifan Mai for creating a HELM evaluation platform for EWoK. This work was supported by the Language Mission of the MIT Quest for Intelligence and by the Open Philantropy Foundation. A.I. is additionally supported by funds from Georgia Tech School of Psychology. J.A. is additionally supported by a Sloan Research Fellowship.
@@ Appendix A Methods
A.1 Dataset generation
EWoK-core-1.0 is the result of 5 randomly-sampled filler-substituted versions of the templates in EWoK. Each version in the script below also serves as a random seed for the dataset. Each dataset includes 1 copy of each template, as specified by the num_fillers parameter (i.e., one unique set of fillers will be sampled for this version). The parameter fix_fillers makes it so that fillers are shared across templates within this version (i.e., agent1 is always filled into by Fatima, if Fatima is the first agent name sampled). Regardless of all of these parameters, variable constraints are always respected and additional fillers are sampled to satisfy constraints (e.g., can_bounce=true).
| Domain | Concepts |
| Social Interactions | help, deceive, hinder, cooperate, compete, evade, seek, chase, learn, teach, respect, insult, flirt, comfort, tease, coerce |
| Social Properties | friendly, hostile, trustworthy, untrustworthy, tolerant, bigoted, boastful, humble, dominant, submissive, shy, extroverted, introverted, confident, warm, cold |
| Social Relations | friend, stranger, enemy, boss, subordinate, colleague, romantic partner, teacher, student, landlord, roommate, tenant, parent, sibling, child |
| Physical Interactions | heat, cool, lift, drop, attract, repel, throw, catch, climb on, push, pull, break, fix, collide with, revolve around, approach, kick, touch, sit on, stand on |
| Physical Dynamics | roll, slide, fall, rise, sink, float, grow, shrink, oscillate, spin, accelerate, slow down |
| Physical Relations | bigger, smaller, occlude, occluded, contain, inside, outside, support, supported, attached, touching, connected, block, trail, hang, tied, on, under, surround, cover |
| Material Dynamics | fold, ripple, pour, stir, flap, splash, droop, drip, hang, pile, trickle, disperse, tap, compress, drape, squeeze, break, rip, wrinkle |
| Material Properties | transparent, opaque, smooth, rough, heavy, light, glossy, matte, fragile, sturdy, cold, warm, soft, hard, elastic, inelastic, bouncy, not bouncy |
| Agent Properties | see, not see, hear, not hear, smell, not smell, taste, not taste, feel, not feel, believe, doubt, imagine, prefer, have no preference, intend, not intend, make an effort, make no effort, make more effort, make less effort, choose |
| Quantitative Properties | more, less, fewer, the same, different, a lot of, a little, none, some, all, enough, not enough, many, few, no, the most, the least, the fewest |
| Spatial Relations | left, right, above, below, north, south, east, west, in front of, behind, close, far, toward, away |
The full list of concepts used in EWoK-core-1.0 templates is provided in Table 3.
A.2 Human data collection
@@
To obtain reliable ratings across the full set of EWoK-core-1.0 items, we collected at least 5 responses per item from a total of N=1,262 participants (591 female, 579 male, 27 other/unknown; median age 36; all US residents who reported English as their first language). Details of the main study are reported in Section 4.4, therefore not repeated here.
Data exclusion
Some online participants may respond randomly and therefore need to be excluded. To evaluate the quality of individual participants’ responses, we computed Pearson R between that participant’s responses and an average of other responses for that domain. Participants whose correlations were <0.3 were excluded.
Validating gold dataset labels using human data
We used human data to validate author-assigned gold labels for items in EWoK-core-1.0. First, we identified all items where the majority human label did not match the gold label (e.g., the plausibility of - was higher than -). Then, we manually inspected those items to determine whether we could trace the discrepancy to an author error or to an error in the item generation pipeline. When possible, the generation pipeline was adjusted to generate the correct version of the item; alternatively, the faulty item was excluded from the dataset.
A.3 Evaluated models
The full set of evaluated models are as follows: gpt2-xl (Radford et al., 2019), phi-1 (Gunasekar et al., 2023), phi-1.5, phi-2 (Li et al., 2023), gemma-2b, gemma-1.1-2b-it, gemma-7b, gemma-1.1-7b-it (Gemma et al., 2024), mpt-7b, mpt-7b-chat, mpt-30b, mpt-30b-chat (MosaicML, 2023), falcon-7b, falcon-7b-instruct, falcon-40b, falcon-40b-instruct (Almazrouei et al., 2023), mistral-7b-v0.1 (Jiang et al., 2023), mixtral-8x7b-v0.1 (Jiang et al., 2024), Meta-Llama-3-8B, and Meta-Llama-3-70B (AI@Meta, 2024). All models were accessed via HuggingFace transformers (Wolf et al., 2020), and all experiments were run on a x GB GPU cluster.
A.4 Text completion experiments
@@ A.5 Prompts
A.6 Mixed effects modeling